EVA Report March 8th

Crew 263 EVA 18 Report 08-03-2022

EVA # 19

Author: Marine Prunier

Purpose of EVA: Replacing the batteries for the atmospheric experiments, Analysis of sulfate deposits using a handheld LIBS (Laser-Induced Breakdown Spectroscopy) analyzer.

Start time: 08:12

End time: 12:34

Narrative: EVA Participants entered the airlock at 08:12. Depressurization ended at 08:17. EVA participants checked the rover batteries. The obtained information is:

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

Spirit: Usage Duration – 157.7 Battery Charge – 100%

Opportunity: Usage Duration – 85.8 Battery Charge – 100%

Curiosity: Usage Duration – 169.9 Battery Charge – 100%

The EVA participants took Opportunity and Curiosity at 08:17 and drove on Entrance Road until Cow Dung Road. Batteries of the Atmospheric Experiments -LOAC, Mega-Ares, Purple Air- were replaced. The old batteries were placed on the rover trunks (08:40).

The participants drove on Cow Dung Road to the southern end of Galileo Road 1104 towards Candor Chasma. From there, the canyon was reached by foot. The EVA participants took about one hour to find the entrance of the canyon. Their GPS went out of battery at the beginning of the EVA, and they took the wrong direction. The correct path to the canyon was finally found and was documented with notes and pictures for a potential other EVA in this area.

Inside the canyon, the participants made several stops to proceed on sketches and pictures. A software issue with the LIBS analyzer prevented the participants to use it on the field, therefore several samples of rocks have been taken for further analysis at the station.

At 11:13, the participants left the point of interest inside the Canyon. At 11:55 they reached the rovers parked on Galileo Road. They arrived back near the Atmospheric Experiments at 12:07 and stopped there for 15 minutes to extract data from Mega-Ares. They drove back to the Hab at 12:27.

Opportunity: Usage Duration – 86.3 Battery Charge – 77%

Curiosity: Usage Duration – 170.3 Battery Charge – 76%

They entered Airlock and at 12:34 the pressurization was over.

Destination: Zone G -final expected point of interest is located at coordinate 4251533.157 N 521043.495 E.

Participants: Marine Prunier (EVA leader), Léa Rouverand, Valentine Bourgeois. HabCom was Cerise Cuny.

Elena Lopez-Contreras was supposed to participate to the EVA, but she woke up feeling tired and decided that it would be safer to stay at the Hab and have some rest.

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road, Galileo Road 1104.

Mode of travel: Rover for battery change and then until the entrance of Candor Chasma, By Foot in Candor Chasma.

Vehicles you will be using (If applicable): Curiosity, Opportunity

Astronomy Report March 8th

Name: Marine Prunier

Crew: 263
Date: 08 Mars 2022


Robotic Telescope Requested (choose one): MLC-RCOS16

Objects to be Imaged this Evening: LDN 1627, M51

Images submitted with this report: /

Problems Encountered: LDN 1627 observation was cancelled on Skynet, M51 observations were taken but most of the pictures are blurry and some of them are not pointed on M51.


Solar Features Observed: /

Images submitted with this report: /

Problems Encountered: /

Journalist Report Match 8th

Sol 15 – A new Era for geology

Author: Nicolas Wattelle

Today was a big day for the three adventurous martionauts that went on EVA. 5 hours was planned for Marine, Léa and Valentine. The longest so far!

To tell you more, they were even free of the usual sport session, in order to be in shape all morning long. And according to their sayings, and also their afternoon naps (not for all of them), it was useful! Especially since they reached the goal of the outing: finding sulphate deposit veins! If, like me, you did not know what it looks like, they brought you a picture of it present in the daily pictures!

They left the Hab an hour before the usual one: at 8 am, they first made some maintenance on our devices around the Hab, and then took two of the rovers (cleverly named Curiosity and Opportunity, two of the real Mars “geologists”), heading to North East. They stopped after half an hour driving, and then, searched for (not without any difficulty) the entrance of a place named Candor Chasma.

The EVA was planned in Candor Chasma because it is a place where we might easily observe Jurassic outcrops. It took place in the second part of the geology field study: now, the crew (to be honest, mainly Léa and Marine who are in charge of it) is focused on documenting the presence of Jurassic sulphate deposit in the region. We are not focused on the transition between the two geological eras anymore. It is good to see this experiment moving forward!

Hence, as planned, it was a sporty one, but it went well. They came back to the Hab with precious samples and a big hunger! Fortunately, Mathéo and I were expecting them with a huge dish of Bolognese pasta that we cooked! They told us the whole story, and we had a great lunch.

The afternoon was calmer than previous ones in the Hab, to allow the ones who needed it to rest a little. The “unlucky” ones had to perform human factors experiments, the others went through data analysis.

The event that woke up the atmosphere a little was a tasting of the famous “crêpes Bonne-Maman” made by Valentine. The crew enjoyed the acrobatic cooking as much as the tasting!

I asked my crewmates around me a new question:

What are your “small pleasures” here during this mission?

For Mathéo, cooking bread makes him proud each time he does it!

Léa appreciates very much eating his freshly cooked bread! Along with Cerise and Valentine, they pointed out the evenings after sending our reports, when we gather to discuss, to watch movies…

And, during our work, we all agreed with Cerise, on the fact that working in front of the desert in the Science Dome is something unique. We will miss seeing Mars through this window!

Operations Report March 8th

SOL: 15

Name of person filing report: Mathéo Fouchet

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report


Spirit rover used: No

Hours: 157.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 86.3

Beginning charge: 100%

Ending charge: 77%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 170.3

Beginning charge: 100%

Ending charge: 76%

Currently charging: Yes

Perseverance rover used: No

Hours: 228.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes and comments: Nothing to report

Summary of Hab operations:

WATER USE: 15 gallons

Water (static tank): 262 gallons

Water (loft tank): 40 gallons

Water Meter: 01563385 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: Nothing to report

Summary of suits and radios: There is a crack under one of the screws present on the red ring of a suit’s helmet. We are doing our best to take care of the material that has been provided to us. I may need help to fix it, I thought about using glue to re-stick both parts together. Let me know if you have any advice on how to fix this issue.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations:

Dual split: Off

SOC: 92% at 6:32pm

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Sol Summary March 8th

Crew 263 Sol Summary 08-03-2022

Sol: 15

Summary Title: Geological field study in Candor Chasma

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: The EVA team left at 8:00 AM for a long EVA in Candor Chasma. They found exactly what they were looking for: satin spar and various forms of sulfate deposits. Sadly, the LIBS had a technical issue (that is impossible to prevent). Even if they took samples, they need more data from this area. They were very impressed by the canyon itself, and mesmerised by the landscape on Galileo Rd. Meanwhile, Nicolas, Mathéo and I did an abs sport session and Elena took a nap.

In the afternoon, the morning EVA team rested while others went through cognitive and teleoperation tasks. We have mixed feelings about the end of the sim being so close. On one hand, some are eager to call their families and friends and go back to the frantic world of communication. On the other hand, the end of the sim marks the end of a one-year-long project, and we kind of already feel nostalgic.

Look Ahead Plan: Tomorrow, 4 crew members will do the same EVA as today’s to collect the desired data. Everyone will perform the ultrasound experiment, as subject and operator.

Anomalies in work: None

Weather: Clear sky.

Crew Physical Status: Very good.

EVA: EVA 19 in the morning. No issue to report except a LIBS malfunction.

Reports to be filed: Sol Summary, EVA Report (n°19), EVA Request (n°20), Operations Report, Journalist Report, HSO Report, Astronomy Report.

Support Requested: No support requested.

Copyright © The Mars Society. All rights reserved. | Main Site